OpenStudy (anonymous):

PLEASE HELPPPP!!!!!!! What is the equation of the line in slope-intercept form that passes through the point (3, –5) and is perpendicular to the line y = 3x – 6? OPTIONS: A) y = –3x – 4 B) y = –3x + 4 C) y= -1/3x - 4 D) y= -1/3x + 4

OpenStudy (anonymous):

Easy, we'll talk through it. Do you know how to find the slope of the perpendicular line?

OpenStudy (anonymous):

is it y=mx+b? or y2-y1/x2-x1??

OpenStudy (anonymous):

Neither for this. The slope of a line is the opposite reciprocal of the slope of the line that is it is perpendicular to.... What is the slope of the equation that you have in the question? (Hint: m is equal to slope)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Right, so in fraction form, that is 3/1... The opposite reciprocal is that fraction, flipped, and with a negative sign in front of it. So what do you have when you do that?

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

Alright awesome :) So we know it has to be either C or D. Now, we can use the formula you wrote.. y=mx+b We know that m (the slope) is -1/3 so we have y=(-1/3)x+b The question gives you a point, so you can plug that point in for x and y. -5=(-1/3)* -3+b. Now, solve for b... Let me know what you got and I'll tell you if that's what I got.

OpenStudy (anonymous):

oops I wrote that wrong... -5=(-1/3)* 3+b...

OpenStudy (anonymous):

okay -5=-1 4

OpenStudy (anonymous):

-5=-1+b Add 1 to both sides so that you have b by itself... -4=b. Plug that into your equation and you have the answer :)

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

@brookew41